OpenStudy (anonymous):

Match each equation with the description of the first step you can take to solve the equation. Match Term Definition k−sixths minus eight−fifteenths equals negative four−fifteenths A) Subtract four−fifteenths from both sides. negative four−fifteenths plus k−sixths equals negative eight−fifteenths B) Subtract eight−fifteenths from both sides. k−sixths plus four−fifteenths equals negative eight−fifteenths C) Add eight−fifteenths to both sides. eight−fifteenths plus k−sixths equals negative four−fifteenths D) Add four−fifteenths to both sides.

OpenStudy (joannablackwelder):

k/6 -8/15=-415 A good first step is to combine like terms while getting the constant term on the opposite side of the equals sign from the x term.. Do you know how to do that?

OpenStudy (anonymous):

no @JoannaBlackwelder

OpenStudy (joannablackwelder):

Add 8/15 to both sides :) That gets the -8/15 on the right to combine with the -4/15 and away from the x term.

OpenStudy (joannablackwelder):

Make sense?

OpenStudy (joannablackwelder):

@izzyoct05 ?

OpenStudy (anonymous):

yes thank you. is that the first questions answer?

OpenStudy (anonymous):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

Yes :)

OpenStudy (joannablackwelder):

Can you try the next one?

OpenStudy (anonymous):

yes :)

OpenStudy (joannablackwelder):

:)

OpenStudy (anonymous):

would i add 4/15 to both sides?

OpenStudy (joannablackwelder):

Perfect! :)

OpenStudy (joannablackwelder):

And the next one?

OpenStudy (joannablackwelder):

I gotta go soon, but I'll be back on later on. :)

OpenStudy (joannablackwelder):

Happy studies and good job so far!

OpenStudy (anonymous):

for the next one would you subract 4/15 from both sides?

OpenStudy (anonymous):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

Yup!

OpenStudy (anonymous):

i got the other ones. thank you! can you help me with some more if i need any help? @joannablackwelder

OpenStudy (joannablackwelder):

Sure. Just tag me. :)

OpenStudy (anonymous):

i have been working on this one for over an hour and i just need the answer and help. Subtract six and one-half minus four and seven-tenths. Justify your steps.

OpenStudy (anonymous):

@joannablackwelder

OpenStudy (joannablackwelder):

Wow, an hour?!

OpenStudy (joannablackwelder):

I would first convert from mixed numbers to improper fractions.

OpenStudy (joannablackwelder):

Multiply the denominator by the whole number and add to the numerator.

OpenStudy (joannablackwelder):

6 1/2 = 13/2

OpenStudy (joannablackwelder):

Make sense, @izzyoct05 ?